Beyond Reform, Conservative, and Orthodox: Aspiring To Be A Serious Jew

Dennis Prager  Ultimate Issues

Summer 1988

(Reprinted without permission)

It is generally believed that Reform, Conservative, and Orthodox define the greatest differences among Jews. This is often untrue, and it is increasingly unconstructive. Jewish life would be immeasurably enriched if instead of focusing on denominational descriptions and goals, Jews would focus on a distinction that is simpler, more accurate, and far more constructive: Serious, and Non-Serious Jews.

These terms are often more accurate than denominational labels because there are serious and non-serious Jews in all of the denominations (though, admittedly, not in equal proportions), and because there are serious Jews who do not fit any denominational label.

But far more important than achieving accuracy, by substituting serious for denominational labels we would achieve two seemingly incompatible goals. Serious makes more demands on Jews, and at the same time it reduces divisiveness among Jews. As more and more Jews would aspire to become serious Jews — as the term will be defined — gratuitous rivalry and sinat khinam (causeless hatred) in Jewish life would decrease while Jewish commitment would increase.

This is not some abstract theory. I have devoted most of the past 20 years to bringing Jews to Judaism, and this non-denominational appeal to Jewish seriousness has been at the core of the effort. When-ever I make the case for Jewish commitment, I make it clear that I am uninterested in whether a Jew becomes Reform, Conservative, or Orthodox, or even falls between denominational cracks. What is demanded of a Jew is that he or she become a serious Jew. This approach has been extra-ordinarily effective.

On a personal note, the approach has enabled me to become one of the very few Jews invited by Reform, Conservative, and Orthodox congregations to lecture on Judaism. And it explains why the book Rabbi Joseph Telushkin and I wrote, \underline{The Nine Questions People Ask About Judaism}, is as widely used to teach Judaism by Chabad (Hasidim) as by Reform Temples.

When asked to become serious Jews rather than, or in addition to, denominational ones, Jews are more likely to begin to take Judaism seriously. Too often, it appears that the greatest interest of Reform, Conservative, and Orthodox Jews is Reform, Conservative, and Orthodox Judaism, not Judaism. This is evident when Orthodox Jews do not rejoice when a previously secular Jew begins to learn Hebrew and pray regularly at Reform services or begins to observe Kashrut thanks to a Conservative congregation. And it is evident when non-Orthodox Jews do not celebrate when an unaffiliated Jew adopts Orthodox Judaism.

The Definition

The serious Jew meets four criteria:

1. This Jew is committed to each of Judaism’s three components: G-d, Torah, and Israel.

2. This Jew attempts to implement the higher ideals of each of these components.

3. Whatever Jewish laws this Jew does or does not observe is the product of a struggle.

4. This Jew is constantly growing in each of these areas.

1 Comittment To G-D, Torah, Israel

Judaism’s three components, G-d, Torah, and Israel (G-d, law, and peoplehood) are so important that together they may be said to constitute Judaism’s trinity. I use the word Trinity precisely. A Jew’s abandonment of one or more of Judaism’s components is a serious a renunciation of normative Judaism as a Christian’s renouncement of any part of the Christian trinity.

To be a serious Jew, one must attempt to be committed equally to Gd, law, and peoplehood. Imbalance toward any one of these has had terrible consequences. Any Jewish movement or group that has not affirmed all three components eventually changed its position, or disappeared, or ostracized itself from the rest of Jewry.

1.1 Overemphasis On G-D

Perhaps the most dramatic example of this was the emphasis by a handful of Jews some 2,000 years ago on G-d and faith. These Jews spoke of faith as sufficient in itself, and dropped the law and people-hood components of Judaism. These Jews founded another religion — Christianity.

A more recent example of this imbalanced emphasis was Reform Judaism, which originally affirmed Judaism’s G-d component and denied that Jews were a people. Reform Jews who continued to uphold this view either eventually assimilated, converted to liberal Christian denominations, or isolated themselves from the rest of Jewry (e.g., the American Council for Judaism). In order to survive, Reform Judaism changed. Today, Reform Judaism affirms Judaism’s peoplehood at least as much as it does G-d and considerably more than it affirms law.

1.2 Overemphasis On Law

At the other end of the Jewish spectrum have been Jews who emphasized the law to the detriment of peoplehood and even G-d. So negatively does normative Judaism look upon such Jews that the Talmud, the very repository of Jewish law, states that the Holy Temple and the Jewish state were destroyed — the greatest tragedy in pre-Holocaust Jewish history — “because Jews adhered to the letter of the law”! Or, as the great Hassidic Rebbe, Menachem Mendl of Kotzk, put it, for some Jews Halacha (Jewish law) has become avodah zorah, (idol worship) — i.e., the law is more revered than G-d.

Judaism without its laws is not Judaism. But a Judaism that worships law is an ugly caricature of itself. Commitment to G-d and to His people must be as evident in a religious Jew’s life as is his commitment to observing the law. That is why the great Orthodox Chief Rabbi of pre- Israel Palestine, Rav Kook, said that the Holy Temple will be rebuilt with ahavat chinam, causeless love. He did not say that it would be rebuilt with stricter legal observance of Kashrut.

1.3 Overemphasis On Peoplehood

Finally, there are many Jews — today, probably most Jews — who are committed only to the peoplehood component of Judaism. Many of these Jews — who call themselves cultural, nationalist, or secular Jews — have made great contributions to Jewish survival. The State of Israel, for example, was founded largely by Jews who dropped G-d and Torah, and dedicated their lives to Jewish peoplehood.

But, such Jews — as committed to the Jewish people as they may be - are as imbalanced in their Jewish commitment as Jews who study the Talmud all day and do little or nothing for the Jewish people.

Even in terms of what these Jews are most attached to — the Jewish people — their imbalanced commitment will eventually be no more constructive than that of Jews who neglect Jewish peoplehood. For most of their children will not even retain their parents’ commitment to Jewish peoplehood. Secular Jews committed to the Jewish people do not produce secular Jews committed to the Jewish people. Their children either stay secular and have little or no commitment to the Jewish people or they become religious, i.e., committed to G-d and Torah, as well as to Israel.

Just as society at large can tolerate, indeed may even need, a tiny number of people fanatically dedicated to art or to science as opposed to all other concerns, Jewish society can tolerate a few Jews who overemphasize G-d or Torah or Israel. But for most of us — for the serious Jew — a deep and equal commitment to each of Judaism’s three components is necessary.

2 The Components Are Not Ends In Themselves

Faith in G-d, observance of law, and identification with Jewish peoplehood are not ends in themselves, however. The serious Jew must attempt to implement the Jewish tasks associated with each component.

2.1 G-d: Ethical Monotheism

It is not enough for a Jew to believe in G-d, or even to love and pray to G-d. The Jewish people has the mission to bring the world to G-d and to his ethics, specifically the Seven Noahide Laws.^{1}

We are here to teach to world the terrible consequences of G-d without ethics (e.g., Crusaders and Khomeinis), and of ethics without G-d (e.g., Nazism and Communism). Unfortunately, however, few Jews of any persuasion take this task seriously. Secular Jews, by definition, do not. Reform Judaism is now so identified with preaching social justice that it is becoming almost indistinguishable from secular humanism. Conservative Judaism is primarily concerned with academic scholarship and commitment to Israel. And Orthodoxy has, by and large, become withdrawn from the world, having become too preoccupied with the Halacha to bother with the world. Bringing the world to G-d and His laws strikes most Orthodox Jews as either naarishkeit (wasteful neglect of the Torah), or pointless until all Jews are brought back to Judaism.

It is a great loss to the Jews and to the world that after the Nazi Holocaust and Communist genocides, Reform, Conservative, Orthodox, and secular Jews continue with these preoccupations as if nothing has changed. It is a world tragedy because Jews are neglecting to teach the Jewish solution to human evil — unlike humanism, Judaism teaches the necessity of G-d for and ethical world; and unlike Christianity, Judaism teaches the primacy of acts over faith. And this neglect is a Jewish tragedy, for every day that Jews do not bring the world closer to G-d and His laws, the Jews are one day closer to another Holocaust.

1These are seven laws that all Judaism holds the children of Noah, i.e., all mankind, responsible to observe. They prohibit idolatry, blasphemy, murder, adultery, stealing, and eating the limb of a living animal, and they demand setting up courts of justice. The Lubavitcher Rebbe, the leader of the Chabad movement, has recently called upon Chabad Hasidim to begin advocating observance of these laws among non-Jews.

2.2 Law: Ethics and Holiness

As regards the law component, the serious Jew acknowledges that just as faith in God is not an end in itself, so to, observance of Jewish law is not an end in itself.

Observance of the law as an end in itself, rather than as a means to the higher goals of ethics and holiness, can actually lead to evil. As the great Torah sage Nachmanides observed, “It is possible to be disgusting with the Torah’s permission.” We are obligated to take Jewish law seriously, but the law has a higher purpose — to show love to other humans and to love G-d.

Therefore, the Jew who is so preoccupied with legal details that he has lost sight of the forest because of all the legal trees does not meet our criteria of a serious Jew. But the Jew who is so preoccupied with forest that he or she has forgotten that it must be composed of Judaism’s unique trees (laws) also does not meet our criteria of a serious Jew.

2.3 Israel: Ahavat Yisrael

The peoplehood component, too, has higher tasks associated with it. It is not enough to identify with the Jewish people or to “feel” very Jewish. The peoplehood component obligates Jews to concern themselves with the welfare of fellow Jews, both in their neighborhood and on the other side of the world. It obligates Jews to realize the difficult task of ahavat yisrael, loving fellow Jews — even those with different religious attitudes and observances.

It is easy for a Jew to cry at films of the Holocaust, to feel pride when Jews win Nobel Prizes, and to cheer at Israeli victories. The serious Jew attends yet another Soviet Jewry rally, visits Israel yet another time, works for Arab and Ethiopian Jews, and tries to hug Jews of other denominations.

3 Commitment As The Product Of Struggle And Knowledge

The serious Jew’s commitment to G-d and to Jewish law must come through struggle.

3.1 Struggle With G-d

The very name of the Jewish people, “Israel,” means “struggle with G-d.” How could Judaism have made it clearer that it deems struggle a necessary element of a Jew’s life?

For many Jews, however, there is little struggle with G-d. Many Jews who believe in G-d believe in a G-d who is more like a celestial butler upon whom they make demands than the Divine law giver who makes demands upon them.

Even observant Jews can fall into an attitude towards G-d that is more a function of habit and communal practice than of struggle. Perhaps that is why the eminent Orthodox Rabbi, Emanuel Rackman, wrote that “G-d created doubt.” After all, G-d certainly could have created us with certainty as to His existence. Yet, G-d chose not to - perhaps because the struggle to come to G-d is as important as the faith itself.

Then there are the secular Jews who refrain from the struggle altogether. “I just cannot believe in G-d” is a cop out. Judaism demands of the atheist Jew that he struggle to come to know G-d. As Eli Weisel has said, “A Jew may love G-d, or a Jew may fight with G-d; but a Jew may not ignore G-d.”

To be a serious Jew, one may not ignore G-d (or pay attention to Him only during crises). The essence of Judaism is to incorporate G-d into one’s daily life (for which we have all the laws “between man and G-d,” erroneously labeled “ritual” laws).

It also means grappling with G-d intellectually. Secular Jews often dismiss the beliefs of many religious people as intellectually shallow. And they are right. But the non-belief of these secular Jews is often equally as shallow. I have rarely encountered a secular Jew who has grappled intellectually with books or with individuals who argue for G-d, or for a G-d oriented life. Whereas it is almost impossible for a religious person to avoid secular challenges to his thinking — television, print media, schooling, advertising: the whole society is one enormous immersion in secularism — the secular individual must seek out religious challenges to his or her thinking. Therefore, the question that secularists so often ask believers — “Do you ever doubt G-d?” can be just as fairly inverted to ask the secularist, “Do you ever doubt your secularism?”

3.2 Struggle With Law

Perhaps the most important characteristic of the serious Jew is struggle with Jewish observance. Put succinctly, this means that Jews who observe Jewish laws must be able to answer the question, “Why do you observer the laws that you observe?” And less- or non- observant Jews must answer the question, “Why don’t you observe what you don’t observe?”

Neither the Orthodox Jew who observes out of habit, nor the Reform whose non-observance is out of habit meets our qualifications for a serious Jew. For example, the serious Jew who keeps Kosher needs to offer reasons for why he keeps kosher. It is not enough to say, “because it says so in the Torah.” Did he ever struggle with the question of why it says so in the Torah? If all he can say is “It says so in the Torah,” in what way is this Jew’s commitment rationally or intellectually superior to that of a Khomeini-type Muslim who keeps the laws because they are in the Koran, or of a fundamentalist Christian who believes that all non-Christians go to hell because he believes it says so in the New Testament?

I grew up in a world that generally practices Kashrut without meaning or reason. So did millions of other Jews in the last hundred years. And it was this absence of any reasons that led the vast majority of them to abandon Kashrut and other Jewish practices.

Conversely, the Jew who does not observe Judaism’s dietary laws needs to explain his non-observance. “I’m Reform, I don’t have to” is not a serious response. Nor is, “It’s an outdated health code.” No serious study of the Jewish dietary laws — even that of non-Jewish scholars — explains Kashrut in terms of health. The reasons for Kashrut overwhelmingly concern ethics, holiness, and Jewish identity, as well as obedience to G-d’s inscrutable will.

Both observance and non-observance, when done out of habit and devoid of reason, betray the lack of struggle on the part of a Jew. That is why the Talmud says, “If someone says ‘I have struggled and found,’ believe him; if he says ‘I have found without a struggle,’ do not believe him.” He is not a serious Jew.

4 One Must Be Moving

Another reason for the importance of struggle in one’s Jewish commitment is that it helps insure growth. The moment one ceases to struggle, it becomes very tempting to stay where one is.

This may be one reason why the Talmud says that “Where a Baal Teshuvah [a returning or repentant] Jews stands, no Tzaddik [saintly] Jew can stand.” Even though the Tzaddik may actually be more observant than the Baal Teshuvah, the returning Jew has struggled - and, therefore, grown — Jewishly. The question that a Jew must ask himself, therefore, is not “How much am I doing as a Jew?” but “How much am I doing as a Jew compared to last month, last year, 10 years ago?”

For all these reasons, study is considered the religious commandment that is equivalent to all the others combined. Jewish movement is impossible without constant Jewish learning. One cannot intuit Judaism; one must study in order to know what is right, what is Jewish. The liberal Jew who equates Judaism with liberal ethics and, therefore, never consults Jewish sources, which as often as not differ with contemporary liberalism, is not a serious Jew. Neither is the Orthodox Jew who only knows enough Judaism in order to say, “It is forbidden.”

These, then, are some of the characteristics of the serious Jew. This Jew is more likely to be found among Orthodox Jews, but he or she is not necessarily Orthodox. Two Jews equally committed to G-d, Torah, and Israel can indeed differ — obviously within certain normative Jewish bounds — on precisely how to translate that commitment into daily life. Indeed, I pray for the day when this is exactly what transpires in Jewish life. The vision of millions of Jews grappling with G-d, Torah, and Israel and debating with one another how best to live all of them, based not on comfort, but on struggle and learning, is truly messianic.

This is copyrighted material and is reproduced here without permission. If Dennis is “speaking to you” here, please write to him requesting a subscription to \underline{Ultimate Issues}.

Dennis Prager Ultimate Issues 10573 Pico Blvd. Los Angeles, CA 90064